# Augment then Smooth: Reconciling Differential Privacy with Certified Robustness

### v1

@@ A.1 Differential Privacy and DPSGD
A.2 DP-PSAC
@@
We used DP-PSAC Xia et al. (2022) in addition to DPSGD to ensure differential privacy while training. Per-sample adaptive clipping (PSAC) is one of a number of approaches that tries to reduce the bias from per-example clipping Bu et al. (2021, 2022c); Esipova et al. (2023), and was motivated for maximizing the signal to noise ratio of gradient updates. It has shown the best performance on several datasets including MNIST, FashionMNIST and CIFAR10. To compare with DPSGD, we incorporated DP-PSAC into our experiments.
@@
DP-PSAC is similar to DPSGD, with the exception that it employs a different clipping method,
@@
Where $g_{i,t}$ denotes the loss gradient for the $i$th sample at iteration $t$. The motivation of this clipping method is that per-example gradients with small norms come from data points on which the model has already converged, and these gradients are often orthogonal to mini-batch gradients. Hence, small norm gradients should not have disproportionally large contributions to the batch gradient as when using per-sample normalization methods Bu et al. (2022c). Compared to such approaches, PSAC reduces the influence of small norm gradients.
@@ B.1 Randomized Smoothing
Previous works have tackled improving the certified robustness of randomized smoothing methods in a variety of ways. The dominant approach for doing so involves modifications to training the base classifier so as to increase robustness and accuracy under Gaussian perturbations. The simplest approach involves adding noise to inputs during training Cohen et al. (2019), while other works utilize regularization Zhai et al. (2020); Li et al. (2019); Jeong and Shin (2020), ensembling Horváth et al. (2021), and adversarial training Salman et al. (2019). While these modifications have been independently studied in the context of improving the certified accuracy of randomized smoothing classifiers, we are the first work integrating these methods with private training through augmentation multiplicity. We provide additional information on two of these training modification methods mentioned in Section 3.
@@ B.1.1 SmoothAdv
One of the most effective methods for improving the performance of randomized smoothing classifiers utilizes adversarial training of the classifier. The method, SmoothAdv proposed in Salman et al. (2019), was motivated by the idea that to improve certified accuracy at a larger certification radius one needs a classifier that is more robust to local perturbations, and the best known method of achieving that is through adversarial training.
@@
$G(x)=(F*\mathcal{N}(0,\sigma^{2}I))(x)=\mathbb{E}_{\delta\sim\mathcal{N}(0,\sigma^{2}I)}[F(x+\delta)].$ (10)
The goal of SmoothAdv is to find a point $\hat{x}$ that maximizes the loss of $G$ in an $l_{2}$ ball around $x$ for the cross entropy loss. They use a projected gradient descent variant to approximately find $\hat{x}$, and define $J(x^{\prime})=l_{\textrm{CE}}(G(x^{\prime}),y)$ to compute
$\Delta_{x^{\prime}}J(x^{\prime})=\Delta_{x^{\prime}}\biggl{(}-\log\mathbb{E}_{\delta\sim\mathcal{N}(0,\sigma^{2}I)}[F(x^{\prime}+\delta)_{y}]\biggr{)}.$ (11)
Since the expectation in Equation 11 is difficult to compute exactly, a Monte Carlo approximation is used by sampling noise $\delta_{1},...,\delta_{m}\sim\mathcal{N}(0,\sigma^{2}I)$ to approximately compute $\Delta_{x^{\prime}}J(x^{\prime})$,
$\Delta_{x^{\prime}}J(x^{\prime})\approx\Delta_{x^{\prime}}\biggl{(}-\log\biggl{(}\frac{1}{m}\sum\limits_{i=1}\limits^{m}[F(x^{\prime}+\delta_{i})_{y}]\biggr{)}\biggr{)}.$ (12)
Finally, $x^{\prime}$ is updated by taking a step in the direction of $\Delta_{x^{\prime}}J(x^{\prime})$, and the final $x^{\prime}$ is used to train the classifier.
B.1.2 MACER
Similar to the Stability training method from Equation 3, MACER Zhai et al. (2020) also modifies the loss for optimization so that the final model has higher certified accuracy at a larger certified radius. In contrast to SmoothAdv, regularizing the model to be more robust in this sense does not require generation of adversarial examples and instead can be optimized directly. The method achieves this by decomposing the error of the smoothed classifier into a classification error term and a robustness error term. The former captures the error from the smoothed classifier misclassifying a given datapoint and the latter captures the error of a certified radius being too small.

### v3

@@ 2 Preliminaries
Certified Robustness.
Adversarial examples are a well-studied phenomenon in ML, in which an input to a model at test-time is perturbed in ways that do not alter its semantics yet cause the model to misclassify the perturbed input (Biggio et al., 2013; Szegedy et al., 2013; Goodfellow et al., 2014). Formally, for a given labeled datapoint $(x,y)$ and classifier $f$, an $(L_{p},\zeta)$-adversary aims to create an adversarial example $x^{\prime}$ such that $\|x^{\prime}-x\|_{p}<\zeta$ and $f(x^{\prime})\neq y$. Despite much research, the most common defense against adversarial examples remains adversarial training (Goodfellow et al., 2014; Zhang et al., 2019). While adversarial training improves robustness to known algorithms for finding adversarial examples, it does not guarantee that a model will be robust to all adversarial examples. This motivates developing certifiable guarantees of robustness which provide a lower bound $r$ on the distance between a correctly classified input and any adversarial example that may be misclassified (Wong & Kolter, 2018; Raghunathan et al., 2018). This lower bound is also known as the certification radius.
Randomized Smoothing.
A popular approach for establishing certified robustness (CR) guarantees is through probabilistic robustness verification which, with high probability, verifies that no adversarial examples exist within a radius $r$ of the original input (Li et al., 2023). The most common method smooths a classifier by averaging the class predictions of $f$ using a smoothing distribution $\mu$ (Lecuyer et al., 2019; Li et al., 2019; Cohen et al., 2019),
[…]
where $\mathbb{I}[a,b]=1{\iff}a=b$ and 00 otherwise (Li et al., 2023). As computing the integral in Equation (1) is intractable, Monte Carlo sampling is used. We denote the approximation of $\hat{g}$ given by Monte Carlo sampling as $g$. Smoothed classifiers are evaluated in terms of their certified accuracy—the fraction of samples correctly classified when certifying robustness at a given radius $r$.
A tight $L_{2}$ radius was obtained by Cohen et al. (2019) when using isotropic Gaussian noise $\mu=\mathcal{N}(x,\sigma^{2}\mathbf{I})$, where $\sigma$ is a hyperparameter that controls a robustness/accuracy tradeoff. In particular, it was proved that for any base classifier $f$, the Gaussian smoothed classifier $g$ is robust around an input $x$ with radius $r=\frac{\sigma}{2}(\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B}))$ where $p_{A}$ and $p_{B}$ denote the probabilities of $c_{A}$ and $c_{B}$, the most and second-most probable classes returned by $g(x)$, and $\Phi^{-1}$ is the inverse of the standard Gaussian CDF. The exact probabilities $p_{A}$ and $p_{B}$ are not needed; one can use lower $\underline{p_{A}}\leq p_{A}$ and upper $\overline{p_{B}}\geq p_{B}$ bounds instead, approximated by Monte Carlo sampling. The output of the smoothed classifier $g(x)$ is approximated by aggregating the predictions of a base classifier $f(x+\eta)$ for $\eta\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$. As a high dimensional standard Gaussian assigns almost no mass near its mean, ensuring that $g(x)$ is accurate at large certification radii $r$ requires $f$ to be accurate on Gaussian perturbed data (Gao et al., 2022).
Related Work.
Previous works combining DP and CR deviate significantly from the standard DPSGD template, and as a result become more difficult to integrate into DP training pipelines.
Phan et al. (2019) introduced Secure-SGD, the first framework aimed at achieving both CR and DP simultaneously. They used a PixelDP approach (Lecuyer et al., 2019) to attain CR and introduced the heterogeneous Gaussian mechanism, which adds heterogeneous noise across elements of the gradient vector.
Phan et al. (2020) introduced StoBatch which employs an autoencoder (Hinton & Zemel, 1993) and a functional DP mechanism, objective perturbation (Zhang et al., 2012), to reconstruct input examples while preserving DP. Subsequently, reconstructed data is used to train a deep neural network, along with adversarial training (Tramèr et al., 2017) to improve robustness.
Tang et al. (2022) proposed perturbing input gradients during training and introduced the multivariate Gaussian mechanism which allows them to achieve the same DP guarantee with less noise added. Their method TransDenoiser follows the architecture of denoised smoothing, adding differentially private noise to a pre-trained classifier.
We note that these prior approaches either add trainable model components increasing the overall complexity, or lack the flexibility to incorporate the latest adversarial training methods.
3 Method
Training ML models to be both differentially private and certifiably robust poses two main challenges. In this section, we describe these challenges and how our DP-CERT framework meets them by effectively improving certified robustness while preserving privacy.
The first challenge surfaces around the use of adversarial training or augmentations with DPSGD. As shown by Cohen et al. (2019), data augmentations used for training can enhance a model’s CR, however, augmented data points could leak private information about the originals. Previous works combining DP and CR have proposed adding noise or adversarial examples during training, but deviate from the standard DPSGD template to address the privacy risks (Phan et al., 2019; 2020; Tang et al., 2022).
The second challenge is that gradient clipping and noise addition in DPSGD harms the convergence rate of training (Chen et al., 2020; Tramèr & Boneh, 2021; Bu et al., 2023a), while restrictive privacy budgets may require stopping training prior to convergence. Robustness on the other hand suffers when models are not converged, as having large gradients at test points makes finding adversarial examples easier (Fawzi et al., 2018). Strategies must be employed to improve the convergence of private training, or regularize gradient size.
We aim to make CR feasible within the standard pipeline of DPSGD, with state-of-the-art convergence and proper accounting for additional privacy risks. Our DP-CERT framework effectively trains with augmented samples while preserving privacy, and integrates advancements in adversarial training and regularizers to enhance certifiable robustness (Salman et al., 2019; Li et al., 2019; Zhai et al., 2020). A schematic of DP-CERT is given in Figure 1 and the training framework in described in Algorithm 2 below. We describe the components in turn.
Augmentation Multiplicity.
The first step in Figure 1 involves training on augmented data with the aim of improving robustness. For each data point $(x_{i},y_{i})$, we obtain $K$ augmented data points $(x_{i}^{j},y_{i})$, where $j\in\{1,...,K\}$ and $x_{i}^{j}$ is the $j$-th augmented version of $x_{i}$. For notational convenience, we use $x_{i}^{0}$ to denote the original data point $x_{i}$. To improve CR, augmentations may be generated adversarially, but since this is computationally expensive, we also consider Gaussian perturbations which, as shown by Cohen et al. (2019), also can enhance a model’s certified robustness. When using Gaussian perturbations instead of adversarial training, we define $x_{i}^{j}=x_{i}+\eta_{j}$, $\eta_{j}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$ for $j\neq 0$.
With respect to the first challenge, an important component of DP-CERT is how we handle the privacy concerns of training with augmented data points. We employ augmentation multiplicity, introduced by De et al. (2022) as a method for incorporating standard augmentation methods into DPSGD training. The method involves averaging the gradients of multiple input augmentations of the same training sample before clipping and noising. Since all downstream impact to the model weights from sample $x_{i}$ is contained in this averaged gradient, clipping it provides a finite sensitivity as required for the sampled Gaussian mechanism used in DPSGD (Mironov et al., 2019), and no additional privacy cost is incurred. We propose to apply a slightly modified version of this method for improving CR with DPSGD, namely, by leveraging augmentation multiplicity to train the network with Gaussian noised input data, akin to classical methods of regularizing models for CR (Cohen et al., 2019).
The model updates can be expressed as follows
$\theta_{t+1}=\theta_{t}-\lambda_{t}\left[\frac{1}{|B_{t}|}\sum_{i\in B_{t}}\textrm{clip}_{C}\left(\frac{1}{K+1}\sum_{j=0}^{K}\nabla_{\theta}L(\theta_{t};x_{i}^{j},y_{i})\right)+\frac{\rho C}{|B_{t}|}\xi\right].$ (2)
$\theta_{t}$ denotes the model parameters at iteration $t$, $\lambda_{t}$ is the learning rate, $B_{t}$ is the $t$’th batch, $C$ is the clipping bound, $K$ is the number of augmentations, $\rho$ is the noise multiplier, $\xi\sim\mathcal{N}(0,\textbf{I})$, and $\nabla_{\theta}L(\theta_{t};x_{i}^{j},y_{i})$ is the gradient of the loss for data point $(x_{i}^{j},y_{i})$. Note that $j$ starts from 0, which means we include the original samples along with the augmented ones in model training. We discovered that including the original samples empirically improves the robust accuracy.
Adversarial Training.
When not using Gaussian augmentations to achieve better certified accuracy, we incorporate adversarial training by deploying existing attacks to create adversarial examples. Specifically, we integrate SmoothAdv (Salman et al., 2019) into private training, which, given original data $(x,y)$, optimizes
$\operatorname*{arg\,max}_{\|x^{\prime}-x\|_{2}\leq\epsilon}\bigg{(}-\log\underset{\eta\sim\mathcal{N}(0,\sigma^{2}I)}{\mathbb{E}}[F_{\theta}(x^{\prime}+\eta)_{y}]\bigg{)},$ (3)
to find an $x^{\prime}$ $\epsilon$-close to $x$ that maximizes the cross entropy between the smoothed classifier $g_{\theta}(x^{\prime})$, and label $y$. Using Monte Carlo sampling, Objective (3) can be optimized by iteratively computing the approximate gradient
$\nabla_{x^{\prime}}\bigg{(}-\log\bigg{(}\frac{1}{K}\sum_{j=1}^{K}F_{\theta}(x^{\prime}+\eta_{j})_{y}\bigg{)}\bigg{)}.$ (4)
Regularization.
To address the second challenge, the second step in Figure 1 adapts stability regularization to private training in order to minimize the distance between the output probability of the original and augmented examples, thereby improving the robustness to input noise. Stability training (Li et al., 2019) adds a smoothed cross-entropy loss as regularization. Inspired by TRADES (Zhang et al., 2019), we instead use the Kullback–Leibler (KL) divergence with a hyperparameter $\gamma$ controlling the strength of the regularization as:
$L_{\text{stab}}(\theta;x_{i},y_{i})=\sum_{j}L(\theta;x_{i}^{j},y_{i})+\gamma D_{\text{KL}}\big{(}F_{\theta}(x_{i})\,\|\,F_{\theta}(x_{i}^{j})\big{)}.$ (5)
Additionally, we propose integrating MACER (Zhai et al., 2020), an alternative training modification to directly optimize the certified accuracy at larger robustness radii without requiring the costly process of adversarial training. MACER achieves this by decomposing the error of a smoothed classifier into a classification error term and a robustness error term, the latter reflecting whether or not the smoothed classifier is able to certify robustness for a given radius. MACER is further described in Appendix B.1.3.
Algorithm 2 DP-CERT Training
0:  Private training set $D=\{(x_{i},y_{i})\mid i\in[N_{\text{prv}}]\}$, loss function $L(\theta_{t};x,y)$, Parameters: learning rate $\lambda_{t}$, noise scale $\rho$, group size $B$, gradient norm bound $C$, number of augmentations $K$.
1:  Initialize ${\theta}_{0}$ randomly
2:  for $t\in[T]$ do
4:     For each $i\in B_{t}$, generate $K$ augmented samples $x_{i}^{j}$Adversarial or Gaussian augmentations}
5:     $L^{\prime}(\theta_{t};x_{i},y_{i})\leftarrow\frac{1}{K+1}\sum_{j=0}^{K}L(\theta_{t};x_{i}^{j},y_{i})+R(\theta_{t},x_{i}^{j})$Add regularizer, average over augmentations}
6:     ${\mathbf{g}}_{t}(x_{i})\leftarrow\nabla_{\theta}L^{\prime}({\theta}_{t};x_{i},y_{i})$Per-sample gradients contain all information from $x_{i}$ and its augmentations}
7:     $\bar{{\mathbf{g}}}_{t}(x_{i})\leftarrow{\mathbf{g}}_{t}(x_{i})/\max\left(1,\frac{\|{\mathbf{g}}_{t}(x_{i})\|_{2}}{C}\right)$Clip gradients}
8:     $\tilde{{\mathbf{g}}}_{t}\leftarrow\frac{1}{|B_{t}|}\left(\sum_{i}\bar{{\mathbf{g}}}_{t}(x_{i})+\mathcal{N}\left(0,\rho^{2}C^{2}{\mathbf{I}}\right)\right)$Add noise to aggregated gradient}
9:      ${\theta}_{t+1}\leftarrow{\theta}_{t}-\lambda_{t}\tilde{{\mathbf{g}}}_{t}$Gradient descent}
10:  end for
11:  Output ${\theta}_{T}$ and compute the overall privacy cost $(\varepsilon,\delta)$ using a privacy accounting method.
Summary of Method.
Table 1: Instantiations of DP-CERT
Method Augmentation Regularization
DP-Gaussian Gaussian None
DP-SmoothAdv Adversarial None
DP-Stability Gaussian Stability
DP-MACER Gaussian MACER
DP-CERT (Figure 1, Algorithm 2) combines augmentation multiplicity, which was originally introduced to improve the accuracy of models trained by DPSGD, with randomized smoothing for inference, and a variety of adversarial augmentation and regularization approaches, which have been shown to improve certified robustness. Comparing the training algorithms 1 and 2 shows that our method follows the template of DPSGD with the following modifications: adversarial or Gaussian perturbations are generated for each data point in the batch; regularization is optionally applied to the loss function; and augmentation multiplicity is used to compute “per-sample” gradients as the average over all augmented data points associated to $x_{i}$. Randomized smoothing at inference time allows certifying robustness, but improved robustness is a result of the training-time modifications.
While DP-CERT is a combination of known methods we emphasize its simplicity compared to previous works that bring CR guarantees to DP. Decreasing complexity is vital for the widespread adoption of robust and private AI, and we simultaneously report improved performance.
Privacy Guarantee.
As described above, augmentation multiplicity achieves the same privacy guarantees as DPSGD by averaging over all augmentations before clipping gradients. This remains true whether random or adversarial augmentations are used, and for any regularized loss function. Once a model is trained with a DP guarantee, it can be used for inference with randomized smoothing at no further privacy cost, thanks to the postprocessing guarantee of DP (Dwork & Roth, 2014; Abadi et al., 2016). Hence, DP-CERT requires no special accounting, and is easily implemented within any mature code package for DP. We note that Gaussian augmentations are a form of input perturbation and may contribute to stronger privacy guarantees. However, input perturbation would first require data points to be clipped for finite sensitivity, impacting performance. Also, we keep the original datapoint that does not have noise added, which is beneficial to performance but not aligned with input perturbation. Hence, we do not attempt to account for DP guarantees from this source of randomness.
In the following sections, we experimentally benchmark four instantiations of the DP-CERT framework as summarized in Table 1.
4 Conclusion
We achieve better certified robustness with DPSGD training through augmentations, regularization, and randomized smoothing, reconciling two crucial objectives for trustworthy ML, namely privacy and robustness. We rely on DP training with augmentations that does not incur additional privacy costs, while employing various regularizations, and adversarial training methods to enhance robustness. Our resulting DP-CERT framework is modular and supports multiple combinations of these methods. Through our extensive experimental study, we confirm that DPSGD training alone, even with state-of-the-art convergence, does not provide satisfactory certified robustness. However, introducing a small number of computationally inexpensive augmentations into training by adding Gaussian noise suffices to yield strong privacy protection and certified robustness, even surpassing much more complex prior methods. By thoroughly analyzing per-sample metrics, we show that the certified radius correlates with the local Lipschitz constant and smoothness of the loss surface; this opens a new path to diagnosing when private models will fail to be robust. In conclusion, our proposed method greatly simplifies the existing solutions to simultaneously achieve CR and DP, and our practical recommendations provide a valuable contribution toward trustworthy ML. When training from scratch, Gaussian augmentations (not adversarial) should be used with DPSGD, and randomized smoothing applied at inference time. For fine-tuning pretrained models, adding stability regularization also helps accuracy, and leads to much lower local Lipschitz constants.
Acknowledgments.
Broader Impact Statement
Nowadays, machine learning finds extensive application in our society. As a result, ensuring the integrity of the models we build is vital. This entails safeguarding individuals’ privacy while maintaining the models’ robustness. Our paper focuses on examining the intersection of privacy and robustness for machine learning models, both of which are essential for establishing trustworthy ML. We introduce the DP-CERT framework, which, compared to prior work, simplifies the process for machine learning developers to create models that are both robust and private, while increasing the robustness afforded for a given privacy level.
While both differential privacy and certified robustness can provide mathematical guarantees on the privacy and robustness of trained models, the guarantees are probabilistic and can randomly fail. For DP, failures roughly occur with probability $\delta$, while for CR, some specific datapoints may not be certifiable for adequately large radii $r$. Practitioners should evaluate their risk tolerances for failures of privacy and robustness, and choose the parameters of these guarantees appropriately.
Furthermore, it has been noted that differential privacy can have particularly negative impacts on the fairness of models (Bagdasaryan et al., 2019; Xu et al., 2021; Esipova et al., 2023). While we do not specifically address fairness aspects in this paper, it is another prong of trustworthy ML that should be taken into account and ensured for any ML model that is to be deployed.
Appendix A Differential Privacy Background
A.1 Differential Privacy and DPSGD
Differential Privacy (DP) (Dwork et al., 2006) is a formal framework that aims to provide mathematical guarantees on the privacy of individual data points in a dataset, while allowing one to learn properties over the entire dataset. More formally a randomized mechanism $M$ fulfills $(\varepsilon,\delta)$-DP if, for all pairs $D$ and $D^{\prime}$ of neighbouring datasets (i.e. datasets differing in only one data point) and all sets $S$ of output values of $M$,
where $\varepsilon$ and $\delta$ are privacy parameters. $\varepsilon$ indicates the level of obtained privacy, with a smaller value of $\varepsilon$ indicating a stronger privacy guarantee. $\delta$ accounts for the probability that the algorithm will suffer a catastrophic privacy failure, which occurs when the privacy loss random variable becomes infinite (Kifer et al., 2022). This will happen when an event with zero probability for one dataset has non-zero probability for a neighbouring dataset, meaning that there is no possible multiplicative factor $e^{\varepsilon}$ that could satisfy the DP inequality Eq. 6 with $\delta=0$. Hence, a larger value of $\delta$ increases the possibility of unmitigated privacy leakage.
The most popular algorithm for implementing DP guarantees in ML is differentially private stochastic gradient descent (DPSGD) (Abadi et al., 2016). It extends the standard SGD algorithm with two additional steps, namely gradient clipping and noise addition. While the former bounds the sensitivity of the model update, the latter implements the privacy guarantee by preventing the gradients from revealing too much information about individual data points. More formally, let $\theta_{t}$ denote the model parameters at training iteration $t$. At each iteration $t$, DPSGD computes the gradient of the loss function with respect to $\theta_{t}$ at an individual data point $x$ as $\nabla_{\theta}L(\theta_{t},x)$. The data point’s gradient is then clipped to a maximum norm $C$ using the operation $\text{clip}\left(\nabla_{\theta}L(\theta_{t},x),C\right)$, which replaces the gradient with a vector of the same direction but smaller magnitude if its norm exceeds $C$. The clipped gradient from all datapoints in a batch are aggregated, then perturbed by adding random noise from $\mathcal{N}(0,\rho^{2}C^{2}\mathbf{I})$, where $\rho$ is the noise scale parameter. We detail DPSGD in Algorithm 1.
A.2 PSAC
@@
Per-sample adaptive clipping (PSAC) (Xia et al., 2023) is one of a number of approaches that tries to reduce the bias from per-example clipping (Bu et al., 2023a; b; Esipova et al., 2023), and was motivated for maximizing the signal to noise ratio of gradient updates. It has shown the best performance for differentially private models on several datasets including MNIST, Fashion-MNIST and CIFAR10. To complement DPSGD as an algorithm with better convergence, we incorporated PSAC into our experiments.
@@
PSAC is similar to DPSGD, with the exception that it employs a different clipping method,
@@
Where $g_{i,t}$ denotes the loss gradient for the $i$th sample at iteration $t$. The motivation of this clipping method is that per-example gradients with small norms come from data points on which the model has already converged, and these gradients are often orthogonal to mini-batch gradients. Hence, small norm gradients should not have disproportionally large contributions to the batch gradient as when using per-sample normalization methods (Bu et al., 2023b). Compared to such approaches, PSAC reduces the influence of small norm gradients, and empirically shows better convergence and utility.
@@ B.1 Randomized Smoothing
Previous works have tackled improving the certified robustness of randomized smoothing methods in a variety of ways. The dominant approach for doing so involves modifications to training the base classifier so as to increase robustness and accuracy under Gaussian perturbations. The simplest approach involves adding noise to inputs during training (Cohen et al., 2019), while other works utilize regularization (Zhai et al., 2020; Li et al., 2019; Jeong & Shin, 2020), ensembling (Horváth et al., 2022), and adversarial training (Salman et al., 2019). While these modifications have been independently studied in the context of improving the certified accuracy of randomized smoothing classifiers, our work is the first to integrate these methods with private training through augmentation multiplicity. We provide additional information on two of these training modification methods mentioned in Section 3, as well as Consistency regularization used in one ablation experiment.
@@ B.1.1 SmoothAdv
One of the most effective methods for improving the performance of a randomized smoothing classifier utilizes adversarial training of the classifier. The method, SmoothAdv proposed in (Salman et al., 2019), was motivated by the idea that to improve certified accuracy at a larger certification radius one needs a classifier that is more robust to local perturbations, and the best known method of achieving that is through adversarial training.
@@
$G(x)=(F*\mathcal{N}(0,\sigma^{2}I))(x)=\mathbb{E}_{\eta\sim\mathcal{N}(0,\sigma^{2}I)}[F(x+\eta)].$ (8)
The goal of SmoothAdv is to find a point $\hat{x}$ that maximizes the cross entropy loss of $G$ in an $l_{2}$ ball around $x$. SmoothAdv uses a projected gradient descent variant to approximately find $\hat{x}$, and define $J(x^{\prime})=L_{\textrm{CE}}(G(x^{\prime}),y)$ to compute
$\nabla_{x^{\prime}}J(x^{\prime})=\nabla_{x^{\prime}}\biggl{(}-\log\mathbb{E}_{\eta\sim\mathcal{N}(0,\sigma^{2}I)}[F(x^{\prime}+\eta)_{y}]\biggr{)}.$ (9)
Since the expectation in Equation 9 is difficult to compute exactly, a Monte Carlo approximation is used by sampling noise $\eta_{1},...,\eta_{m}\sim\mathcal{N}(0,\sigma^{2}I)$ to approximately compute $\nabla_{x^{\prime}}J(x^{\prime})$,
$\nabla_{x^{\prime}}J(x^{\prime})\approx\nabla_{x^{\prime}}\biggl{(}-\log\biggl{(}\frac{1}{m}\sum\limits_{i=1}\limits^{m}[F(x^{\prime}+\eta_{i})_{y}]\biggr{)}\biggr{)}.$ (10)
Finally, $x^{\prime}$ is updated by taking a step in the direction of $\nabla_{x^{\prime}}J(x^{\prime})$, and the final $x^{\prime}$ is used to train the classifier.
B.1.2 Consistency Regularization
For ablation purposes, we also compare to Consistency regularization (Jeong & Shin, 2020), a method very similar to Stability training which instead minimizes the KL divergence between $\hat{F}_{\theta}(x_{i})$ and $F_{\theta}(x_{i})$, where $\hat{F}_{\theta}(x_{i})=\frac{1}{K}\sum_{j}F_{\theta}(x_{i}^{j})$ is the average output probability of all smoothed samples. The loss can be expressed as
$\ L_{\text{cons}}(x_{i},y_{i})=\sum_{j}L_{\textrm{CE}}(x_{i}^{j},y_{i})+\gamma D_{\text{KL}}\big{(}\hat{F}_{\theta}(x_{i})||F_{\theta}(x_{i}^{j})\big{)}.$ (11)
B.1.3 MACER
Similar to the Stability training method, MACER (Zhai et al., 2020) also modifies the loss for optimization so that the final model has higher certified accuracy at a larger certified radius. In contrast to SmoothAdv, regularizing the model to be more robust in this sense does not require generation of adversarial examples and instead can be optimized directly. The method achieves this by decomposing the error of the smoothed classifier into a classification error term and a robustness error term. The former captures the error from the smoothed classifier misclassifying a given datapoint and the latter captures the error of a certified radius being too small.